Following information is available for Companies Ace Ltd. and Pace Ltd.: (₹ in lacs)

Particulars Ace Ltd. Pace Ltd. Long term Debt 625 700 Equity 2100 2850 Current assets 450 550 Current liabilities 300 375 Net Profit 115 178 Revenue (net) 355 452

a. Compute Debt-equity ratio, current ratio for both companies.


1
Expert's answer
2022-03-17T09:24:02-0400

Debt Equity Ratio = (Short term Debt + Long Term Debt) / Shareholders Equity 

Ace Ltd 

Debt Equity Ratio = (625+300)/2100

Debt Equity Ratio =925/2100

Debt Equity Ratio = 0.44

Pace Ltd

Debt Equity Ratio = (700+375)/2850

Debt Equity Ratio = 1075/2850

Debt Equity Ratio = 0.38

Current Ratio = Current Assets/Current Liabilities 

Ace Limited  Current Ratio = 450/300 Current Ratio = 1.5

Pace Limited  Current Ratio = 550/375 Current Ratio = 1.5